OpenStudy (anonymous):

I have a problem here and made it half way (not sure how to do the rest) (2-i)(-3+6i) FOIL I got: -6 + 12 i +3 i - 6 i^2

OpenStudy (primeralph):

What is i^2 ?

OpenStudy (anonymous):

i squared

OpenStudy (whpalmer4):

Are you multiplying complex numbers, such that \(i= \sqrt{-1}\)? If so, what does \(i^2=\)?

OpenStudy (anonymous):

yes and i squared is -1

OpenStudy (anonymous):

im not sure how I would combine like terms then plug in -1 for i squared

OpenStudy (whpalmer4):

Right, so you can do a bit more simplification of your result. Combine the \(i\) terms, and change the \(i^2\) to \(-1\).

OpenStudy (whpalmer4):

\[(2-i)(-3+6i) = 2*(-3)+2*6i-i*(-3) -i*6i = -6+12i + 3i -6i^2\]\[=-6+15i-6i^2\]Right?

OpenStudy (whpalmer4):

Now substitute \(-1\) for \(i^2\): \[-6+15i-6*(-1) = -6 + 15i + 6 = \]

OpenStudy (anonymous):

(2-i)(-3+6i) -6+12i+3i-6i^2 is when you foil. -6+15i+6 the -6 and +6 cancel so you are left with your answer of; 15i

OpenStudy (anonymous):

yes I see so the 6's cancel out

OpenStudy (anonymous):

Yes (:

OpenStudy (anonymous):

thanks so much!!! to you both